# Data-adaptive structural change-point detection via isolation

### v2

@@ 2.3 Location of the largest difference and the isolation aspect
Note that the lower bound of $n$ ensures that the right hand side of the second inequality in (6) is non-negative. The result in (6) both guarantees detection in an interval where the change-point is isolated but also allows for the detection of the neighbouring change-point. This is because the end-point of the interval the algorithm will be applied to after detection of the first change-point (either $r_{J}$ or $r_{J+1}$), will have a distance of at least $\delta_{T}/2n$ from all the undetected change-points; see (C) and (D) in the Appendix for a proof that $\delta_{T}/2n$ is the minimum distance of the end-points from the change-point that guarantees detection.
It should be mentioned that $\tilde{\delta}_{j}=\mathcal{O}(\delta_{T})$ for any $j=0,1,\ldots,N$. Also, since $2n\lambda_{T}\leq\delta_{T}$, $n\lambda_{T}=\mathcal{O}(\delta_{T})$ and, from (6), the upper bound of $\delta_{s,e}^{J}$ is $\mathcal{O}(\delta_{T})$. Now, as is given by assumption (A2) of Section 3.1, it must hold that $\sqrt{\delta_{T}}\underline{f}_{T}\geq\underline{C}\sqrt{\log T}$, where $\underline{f}_{T}=\min_{j}|f_{r_{j}+1}-f_{r_{j}}|$ is defined as the minimum magnitude of jumps. This assumption implies that when $T\rightarrow\infty$ then $\sqrt{\delta_{T}}\underline{f}_{T}$ increases. If $\delta_{T}$ increases then the upper bound of $\delta_{s,e}^{J}$ increases with the same rate. If $\underline{f}_{T}$ increases, the signal-to-noise ratio becomes larger and so the probability that the largest difference is exactly at the location of the change-point increases. More details on this, in the case that $\epsilon_{t}$ is Gaussian, are provided in Section 3.3. In either case, the probability that $d_{s,e}$ lies within $\delta_{s,e}^{J}$ of any change-point increases as $T$ becomes large. We can reach the same conclusions using Assumptions (A3) and (A4) for the case of univariate piecewise-linear signals and multivariate signals, respectively.
Considering now any $d_{s,e}\in\{1,2,\ldots,T\}$ and using (6), it must hold that
@@
$d_{s,e}\in\left[1,r_{1}+\frac{\tilde{\delta}_{1}}{2}-\frac{3\delta_{T}}{4n}\right]\cup\left[r_{N}-\frac{\tilde{\delta}_{N-1}}{2}+\frac{3\delta_{T}}{4n},T\right]\cup\left\{\bigcup_{j=2}^{N-1}B_{n;j}\right\}$ (7)
@@
$B_{n;j}=\left[r_{j}-\frac{\tilde{\delta}_{j-1}}{2}+\frac{3\delta_{T}}{4n},r_{j}+\frac{\tilde{\delta}_{j}}{2}-\frac{3\delta_{T}}{4n}\right].$ (8)
Note that the length of the interval where $d_{s,e}$ cannot lie under this assumption is $2(\lceil 3\delta_{T}/4n\rceil-1)$ around the midpoint of two consecutive change-points. This can be as small as $23\lambda_{T}-$ 2, which means that for appropriately chosen values of $n$ and $\lambda_{T}$, any location is an acceptable location for $d_{s,e}$ besides the midpoint between consecutive change-points.
@@ 3 Theory
DAIS requires knowledge of the standard deviation, $\sigma$, of the data generation mechanism in (1). If $\sigma$ is unknown, it can be estimated using the Median Absolute Deviation method (MAD), with the estimator defined as $\hat{\sigma}:=C\times\text{median}(\left|\boldsymbol{x}-\text{median}(\boldsymbol{x})\right|)$ for $\boldsymbol{x}=(x_{1},x_{2},\dots x_{T})$, as proposed by [26]. In the case of Gaussian data, for $C=1.4826$, $\hat{\sigma}$ is a consistent estimator of the population standard deviation $\sigma$ as proven by [43]. In this section, we assume that $\epsilon_{t}$ in (1) are i.i.d. random variables following the $\mathcal{N}(0,1)$ distribution. This assumption is commonly used in the literature to prove theoretical consistency results. Without loss of generality, we let $\sigma=1$ in order to make notation easier in the proof of the theorems presented in this section. The model becomes
$X_{t}=f_{t}+\epsilon_{t},\quad t=1,2,\ldots,T.$ (9)
In Sections 3.1 and 3.2 we provide the theoretical results on the consistency of the algorithm for the cases of piecewise-constant and continuous, piecewise-linear signals, respectively. The proofs of Theorems 1 and 2 can be found in Appendix C and D, respectively. In Section 6.3 we also provide theoretical results for the consistency in the case of multivariate signals and the outline of the proof can be found in Appendix E. In Section 3.3, we return to the discussion of Section 2.3 and present some results in the case when (7) holds, and $\epsilon_{t}$ in (1) are i.i.d. from the $\mathcal{N}(0,1)$ distribution.
@@ 3.1 Piecewise-constant signals
Under piecewise-constancy, $f_{t}=\mu_{j}$ for $t=r_{j-1}+1,\ldots,r_{j}$ and $f_{r_{j}}\neq f_{r_{j}+1}$. The statistic used as a contrast function in this case is the absolute value of the CUSUM statistic, which is defined as
@@
    The minimum distance, $\delta_{T}$, between two successive change-points and the minimum magnitude of jumps $\underline{f}_{T}=\min_{j}\left|f_{r_{j}+1}-f_{r_{j}}\right|$ are connected by $\sqrt{\delta_{T}}\underline{f}_{T}\geq\underline{C}\sqrt{\log T}$ for a large enough constant $\underline{C}$.
The first assumption is a formal statement of what is discussed in Section 2.3 and is required to ensure that every change-point can be detected in an interval in which it is isolated. It is not a strict assumption as for appropriately chosen values of $n$ and $\lambda_{T}$, (7) includes any point besides the midpoint of two consecutive change-points. The exact probability of the consecutive pairwise difference being larger in absolute value at a change-point compared to a point where no change occurs, in the case where $\epsilon_{t}$ is assumed to be Gaussian, is provided in Section 3.3. (A2) is not a very restrictive assumption, for example if $\delta_{T}\geq O(\log T)$, then $\underline{f}_{T}$ is allowed to decrease as $T$ increases. Further discussion for (A2) is given after Theorem 1.
Below, we provide the relevant theorem for the consistency of DAIS in accurately estimating the true number and the locations of the change-points in the case of piecewise-constant signals.
@@
Let $\{X_{t}\}_{t=1,2,\ldots,T}$ follow model (9), with $f_{t}$ being a piecewise-constant signal and assume that the random sequence $\{\epsilon_{t}\}_{t=1,2,\ldots,T}$ is independent and identically distributed (i.i.d.) from the normal distribution with mean zero and variance one and also that assumptions (A1) and (A2) hold. Let $N$ and $r_{j},j=1,2,\ldots N$ be the number and location of the change-points, while $\hat{N}$ and $\hat{r}_{j},j=1,2,\ldots\hat{N}$ their estimates, sorted in increasing order. In addition, $\Delta^{f}_{j}=\left|f_{r_{j}+1}-f_{r_{j}}\right|,j=1,2,\ldots N$ is the magnitude of each discontinuity in $f_{t}$, $\underline{f}_{T}=\min_{j}\Delta^{f}_{j}$ and $\delta_{T}=\min_{j=1,2,\ldots,N+1}\lvert r_{j}-r_{j-1}\rvert$, where $r_{0}=0,r_{N+1}=T$. Then there exist positive constants $C_{1},C_{2},C_{3},C_{4}$ which do not depend on $T$ such that for $C_{1}\sqrt{\log T}\leq\zeta_{T}<C_{2}\sqrt{\delta_{T}}\underline{f}_{T}$ and for sufficiently large $T$, we obtain
@@ 3.2 Continuous, piecewise-linear signals
For the case of continuous, piecewise-linear signals, $f_{t}=\mu_{j,1}+\mu_{j,2}t$ for $t=r_{j-1}+1,\ldots,r_{j}$ and $f_{r_{j}-1}+f_{r_{j}+1}\neq 2f_{r_{j}}$ with the additional constraint $\mu_{j,1}+\mu_{j,2}r_{j}=\mu_{j+1,1}+\mu_{j+1,2}r_{j}$ so that the signal is continuous. Through a log-likelihood approach, [6] shows that contrast function is $C_{s,e}^{b}\left(\boldsymbol{X}\right)=\left|\langle\boldsymbol{X},\boldsymbol{\phi^{b}_{s,e}}\rangle\right|$, where the contrast vector $\boldsymbol{\phi^{b}_{s,e}}=(\phi^{b}_{s,e}(1),\dots,\phi^{b}_{s,e}(T))$ is given by
@@
Before presenting the theoretical result for the consistency of the number and location of the estimated change-points by DAIS in the case of continuous piecewise-linear signals, we require the following assumption, which is equivalent to (A2) in Section 3.1, but now under the current setting of changes in the slope.
@@
    The minimum distance, $\delta_{T}$, between two successive change-points and the minimum magnitude of jumps $\underline{f}_{T}=\min_{j=1,2,\ldots,N}\lvert 2f_{r_{j}}-f_{r_{j}+1}-f_{r_{j}-1}\rvert$ are connected by $\delta_{T}^{3/2}\underline{f}_{T}\geq C^{\ast}\sqrt{\log T}$ for a large enough constant $C^{\ast}$.
@@
Let $\{X_{t}\}_{t=1,2,\ldots,T}$ follow model (9), with $f_{t}$ being a continuous piecewise-linear signal and assume that the random sequence $\{\epsilon_{t}\}_{t=1,2,\ldots,T}$ is independent and identically distributed (i.i.d.) from the normal distribution with mean zero and variance one and also that (A1) and (A3) hold. Let $N$ and $r_{j},j=1,2,\ldots N$ be the number and location of the change-points, while $\hat{N}$ and $\hat{r}_{j},j=1,2,\ldots\hat{N}$ are their estimates, sorted in increasing order. In addition, $\Delta^{f}_{j}=\left|2f_{r_{j}}-f_{r_{j}+1}-f_{r_{j}-1}\right|,j=1,2,\ldots N$ is the magnitude of each change of slope in $f_{t}$, $\underline{f}_{T}=\min_{j}\Delta^{f}_{j}$ and $\delta_{T}=\min_{j=1,2,\ldots,N+1}\lvert r_{j}-r_{j-1}\rvert$, where $r_{0}=0,r_{N+1}=T$. Then there exist positive constants $C_{1},C_{2},C_{3},C_{4}$ which do not depend on $T$ such that for $C_{1}\sqrt{\log T}\leq\zeta_{T}<C_{2}\delta_{T}^{3/2}\underline{f}_{T}$ and for sufficiently large $T$, we obtain
@@
The proof of Theorem 2 can be found in Appendix D. Again, the lower bound of the probability is $1-\mathcal{O}(1/T)$ and the difficulty of the problem, $\delta_{T}^{3/2}\underline{f}_{T}$ is analogous to $\sqrt{\delta_{T}}\underline{f}_{T}$, which appears in (A2) for the case of piecewise-constant signals. For $\underline{f}_{T}\sim T^{1/2}$, the change-point detection accuracy is $O\left(T^{-1/3}\left(\log T\right)^{1/3}\right)$, which differs from the optimal rate $O(T^{-1/3})$ derived by [41] only by the logarithmic factor. This means that DAIS’s change-point detection rate is almost optimal.

### v3

@@ 2.3 Location of the largest difference and the isolation aspect
guarantees detection in an interval where the change-point is isolated and also allows for the detection of the neighboring change-points. This is because the end-point of the interval the algorithm will be applied to after detection of the first change-point (either $r_{J}$ or $r_{J+1}$), will have a distance of at least $\delta_{T}/2n$ from all the undetected change-points; see (B) and (S3) in the supplementary material for a proof that $\delta_{T}/2n$ is the minimum distance of the end-points from the change-point that guarantees detection. Note that the lower bound of $n$ ensures that the right hand side of the second inequality in (6) is non-negative. As is explained later in detail, for any $n>3/2$, the upper bound in (6) increases as $T\rightarrow\infty$. In the case that $n=3/2$ and $\tilde{\delta}_{J}=\delta_{T}$, the upper bound becomes 00, which means that $d_{s,e}$ has to lie exactly at the location of the change-point for (6) to be satisfied. However, this does not mean that the method fails in this case; it only implies that the change-point is not guaranteed to be isolated when it is detected. More discussion about this is given after Assumption (A1) in Section 3.1.
It is worth mentioning that (6) is derived such that all change-points can be detected in an interval in which they are isolated, for any $\lambda_{T}\leq\delta_{T}/2n$. If we fix $\lambda_{T}=1$, then in the worst-case scenario, which is for detection to happen when both end-points are necessarily at least at a distance $\delta_{T}/2n$ from the true change-point location, one of the end-points will be exactly at a distance $\lceil\delta_{T}/2n\rceil$. So, for $\lambda_{T}=1$, (6) can be relaxed to
$0\leq\delta_{s,e}^{J}\leq\frac{\tilde{\delta}_{J}}{2}-\Big{\lceil}\frac{\delta_{T}}{2n}\Big{\rceil}.$ (7)
For the choice $n=3/2$ and $\tilde{\delta}_{J}=\delta_{T}$, using (7), $d_{s,e}$ has to be within $\delta_{T}/2-\Big{\lceil}\delta_{T}/3\Big{\rceil}$ of the true location of the change-point.
From the consistency result as in Theorem 1, $\delta_{T}$ must be at least of order $O(\log T)$. This is because $\delta_{T}$ needs to be larger than the distance between the estimated and true locations of the change-points, which is at most of order $O(\log T)$, as can be seen from (16); otherwise we would not be able to match the estimated change-points with the true ones. Also, by definition, $\tilde{\delta}_{j}$ is at least of order $\mathcal{O}(\delta_{T})$ for any $j=0,1,\ldots,N$. This means that as $T\rightarrow\infty$, the upper bound of $\delta_{s,e}^{J}$ as given in (7) (and (6) for $n>3/2$) increases. Additionally, if $\underline{f}_{T}$ increases, the signal-to-noise ratio becomes larger and so the probability that the largest difference is exactly at the location of the change-point increases. In both cases, the probability that $d_{s,e}$ lies within $\delta_{s,e}^{J}$ of any change-point increases. More details on this, in the case that $\epsilon_{t},t=1,2,\ldots,T$ follow the Gaussian distribution, are provided in Section 3.3.
We define
$d_{s,e}^{\max}=\left\{\begin{array}[]{cc}T-1,&\textrm{when }f_{t}\textrm{ is piecewise-constant,}\\ T-2,&\textrm{when }f_{t}\textrm{ is continuous, piecewise-linear,}\end{array}\right.,$ (8)
to be the maximum value that $d_{s,e}$ can obtain. In the case of continuous, piecewise polynomial signals of order $p$, it can be extended to $d_{s,e}^{\max}=T-p-1$. Considering now any $d_{s,e}\in\{1,2,\ldots,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}d_{s,e}^{\max}}\}$ and using (6), in order to have detection while the change-point is isolated, it must hold that
@@
$d_{s,e}\in\left[1,r_{1}+\frac{\tilde{\delta}_{1}}{2}-\frac{3\delta_{T}}{4n}\right]\cup\left[r_{N}-\frac{\tilde{\delta}_{N-1}}{2}+\frac{3\delta_{T}}{4n},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}d_{s,e}^{\max}}\right]\cup\left\{\bigcup_{j=2}^{N-1}B_{n;j}\right\}$ (9)
@@
$B_{n;j}=\left[r_{j}-\frac{\tilde{\delta}_{j-1}}{2}+\frac{3\delta_{T}}{4n},r_{j}+\frac{\tilde{\delta}_{j}}{2}-\frac{3\delta_{T}}{4n}\right]$ (10)
and the union over $j$ in (9) is the empty set for $N\leq 2$. Note that the length of the interval where $d_{s,e}$ cannot lie under this assumption, at each step of the algorithm, is $12\lceil 3\delta_{T}/4n\rceil-$ 1 around the midpoint of every pair of consecutive change-points. Similar intervals to (9) can be defined using (7), but we focus on the general $\lambda_{T}\leq\delta_{T}/2n$ case.
Now, let us consider the case where the signal-to-noise ratio is so small that it is impossible for the location of the largest difference to be informative (near anyone of the true change-points); in such cases, and based on the noise, the largest difference could be anywhere in the set $\{1,2,\ldots d_{s,e}^{{\rm max}}\}$, with $d_{s,e}^{{\rm max}}$ as in (8). Assumptions (A2) or (A3), which can be found in Sections  3.1 and 3.2, respectively, and are needed for detection of the change-points to be possible, are considered to be satisfied. In this case, the probability that the location of the largest difference does not lie in (9) is $(N-1)\left(2\lceil\frac{3\delta_{T}}{4n}\rceil-1\right)/d_{s,e}^{\max}$. Therefore, even in the worst case that the signal-to-noise ratio is such that the location of the largest difference is uninformative, as long as $N\delta_{T}$ is of order less that $O(T)$, the probability of the result in (9) will still go to 1 as $T\to\infty$. This is a strong indication that the requirement expressed in (9), which is, in fact, Assumption (A1) provided in Section 3, is not restrictive and it only used to have guaranteed isolation and detection at the same time.
@@ 3 Theory
DAIS requires knowledge of the standard deviation, $\sigma$, of the data generation mechanism in (1). If $\sigma$ is unknown, it can be estimated using the Median Absolute Deviation method (MAD), with the estimator defined as $\hat{\sigma}:=1.4826\times\text{median}(\left|\boldsymbol{x}-\text{median}(\boldsymbol{x})\right|)$ for $\boldsymbol{x}=(x_{1},x_{2},\dots x_{T})$, as proposed by [30]. The obtained estimator has a simple explicit formula, needs little computation time, is very robust with a bounded influence function and a 50% breakdown point, and, in the case of Gaussianity is a consistent estimator of the population standard deviation ([53]). Due to the expected presence of change-points in the data, we use this estimator on the first and second order differenced data, for piecewise-constant and continuous, piecewise-linear signals, respectively; see for example [7] for a similar approach. The exact expressions for the piecewise-constant and continuous, piecewise-linear cases are provided in Sections 3.1 and 3.2, respectively. We take $\sigma$ to be known and, without loss of generality, we let $\sigma=1$. The model becomes
$X_{t}=f_{t}+\epsilon_{t},\quad t=1,2,\ldots,T,$ (11)
where $\epsilon_{t},t=1,2,\ldots,T$ are i.i.d. random variables following the $\mathcal{N}(0,1)$ distribution. This assumption is commonly used in the literature to prove theoretical consistency results.
Before proceeding, we introduce some notation that is used throughout the paper. With $f_{t}$ the underlying signal as in (1), then for $j=1,2,\ldots,N$, the minimum magnitude of change in the data sequence is defined as
$\underline{f}_{T}=\min_{j}\Delta^{f}_{j},$ (12)
where
$\Delta^{f}_{j}=\left\{\begin{array}[]{ll}\left|f_{r_{j}+1}-f_{r_{j}}\right|,&\textrm{when }f_{t}\textrm{ is piecewise-constant,}\\ \\ \lvert 2f_{r_{j}}-f_{r_{j}+1}-f_{r_{j}-1}\rvert,&\textrm{when }f_{t}\textrm{ is continuous, piecewise-linear.}\end{array}\right.$ (13)
Furthermore, for $r_{0}=0$ and $r_{N+1}=T$, then the minimum distance between consecutive change-points is defined as
$\delta_{T}=\min_{j=0,\ldots,N}\tilde{\delta}_{j},\textrm{ where }\tilde{\delta}_{j}=r_{j+1}-r_{j}.$ (14)
In Sections 3.1 and 3.2 we provide the theoretical results on the consistency of the algorithm for the cases of piecewise-constant and continuous, piecewise-linear signals, respectively. The proofs of Theorems 1 and 2 can be found in Appendix B and the supplementary material, respectively. In Section 6.3 we also provide theoretical results for the consistency in the case of multivariate signals and the outline of the proof can be found in Appendix C. In Section 3.3, we return to the discussion of Section 2.3 and present some results in the case when (9) holds, and $\epsilon_{t}$ in (1) are i.i.d. from the $\mathcal{N}(0,1)$ distribution.
@@ 3.1 Piecewise-constant signals
Under piecewise-constancy, $f_{t}=\mu_{j}$ for $t=r_{j-1}+1,\ldots,r_{j}$ and $f_{r_{j}}\neq f_{r_{j}+1}$. The estimator of the standard deviation used is $\hat{\sigma}_{C}:=1.4826\times\text{median}(\left|\boldsymbol{Y}-\text{median}(\boldsymbol{Y})\right|)/\sqrt{2}$, where $\boldsymbol{Y}=(X_{2}-X_{1},\ldots,X_{T}-X_{T-1})$. The statistic used as a contrast function in this case is the absolute value of the CUSUM statistic, which is defined as
@@
    The minimum distance, $\delta_{T}$, between two successive change-points, and the minimum magnitude of jumps $\underline{f}_{T}$, as defined in (14) and (12), respectively, are connected by $\sqrt{\delta_{T}}\underline{f}_{T}\geq\underline{C}\sqrt{\log T}$ for a large enough constant $\underline{C}$.
The first assumption is a formal statement of what is discussed in Section 2.3 and is required to ensure that every change-point can be detected in an interval where isolation is guaranteed. It is not a strict assumption, for more details on this see Sections 2.3 and 3.3. The exact probability of the consecutive pairwise difference being larger in absolute value at a change-point location compared to a point where no change occurs, in the case where $\epsilon_{t}$ are assumed to be Gaussian, is provided in Section 3.3. (A2) is a typical assumption in the literature and characterizes the difficulty of the detection problem. It is worth mentioning that if $\delta_{T}$ is of order greater than $O(\log T)$, then $\underline{f}_{T}$ is allowed to decrease as $T$ increases. Further discussion for (A2) is given after Theorem 1. Assumptions (A1) and (A2), together, guarantee that all change-points will be detected in intervals in which they are isolated. However, even when (A1) fails, the change-points can still potentially get detected in practice. This can happen in one of two ways. The first is that detection occurs in an interval in which the change-point is isolated, whose end-points (one or both) are at a distance of less than $\delta_{T}/2n$ from the true location of that change-point. The second way is in an interval in which the change-point is not isolated. In both cases the detection power, which is based on the value of the contrast function, is reduced, but change-points can still be detected. The detection power is maximized when the change-point is isolated, in the midpoint of the interval under consideration with end-points far from it, and this is what we attempt to achieve using Assumption (A1).
Below, we provide the relevant theorem for the consistency of DAIS in accurately estimating the true number and the locations of the change-points in the case of piecewise-constant signals. The proof of the theorem can be found in Appendix B.
@@
Let $\{X_{t}\}_{t=1,2,\ldots,T}$ follow model (11), with $f_{t}$ being a piecewise-constant signal and assume that the random sequence $\{\epsilon_{t}\}_{t=1,2,\ldots,T}$ is independent and identically distributed (i.i.d.) from the normal distribution with mean zero and variance one and also that Assumptions (A1) and (A2) hold. Let $N$ and $r_{j},j=1,2,\ldots N$ be the number and location of the change-points, while $\hat{N}$ and $\hat{r}_{j},j=1,2,\ldots\hat{N}$ their estimates, sorted in increasing order. Then, with $\Delta_{j}^{f}$, $\underline{f}_{T}$, and $\delta_{T}$ as in (13), (12), and (14), respectively, there exist positive constants $C_{1},C_{2},C_{3},C_{4}$ independent of $T$ such that for $C_{1}\sqrt{\log T}\leq\zeta_{T}<C_{2}\sqrt{\delta_{T}}\underline{f}_{T}$ and for sufficiently large $T$, we obtain
@@ 3.2 Continuous, piecewise-linear signals
For the case of continuous, piecewise-linear signals, $f_{t}=\mu_{j,1}+\mu_{j,2}t$ for $t=r_{j-1}+1,\ldots,r_{j}$ and $f_{r_{j}-1}+f_{r_{j}+1}\neq 2f_{r_{j}}$ with the additional constraint $\mu_{j,1}+\mu_{j,2}r_{j}=\mu_{j+1,1}+\mu_{j+1,2}r_{j}$ so that the signal is continuous. The estimator of the standard deviation used is $\hat{\sigma}_{L}:=1.4826\times\text{median}(\left|\boldsymbol{Y}-\text{median}(\boldsymbol{Y})\right|)/\sqrt{6}$, where $\boldsymbol{Y}=(X_{1}-2X_{2}+X_{3},\ldots,X_{T-2}-2X_{T-1}+X_{T})$. Through a log-likelihood approach, [7] shows that contrast function is $C_{s,e}^{b}\left(\boldsymbol{X}\right)=\left|\langle\boldsymbol{X},\boldsymbol{\phi^{b}_{s,e}}\rangle\right|$, where the contrast vector $\boldsymbol{\phi^{b}_{s,e}}=(\phi^{b}_{s,e}(1),\dots,\phi^{b}_{s,e}(T))$ is given by
@@
Before presenting the theoretical result for the consistency of the number and location of the estimated change-points by DAIS in the case of continuous, piecewise-linear signals, we require the following assumption, which is equivalent to (A2) in Section 3.1, but now under the current setting of changes in the slope.
@@
    The minimum distance, $\delta_{T}$, between two successive change-points and the minimum magnitude of jumps $\underline{f}_{T}$, as defined in (14) and (12), respecitvely, are connected by $\delta_{T}^{3/2}\underline{f}_{T}\geq C^{\ast}\sqrt{\log T}$ for a large enough constant $C^{\ast}$.
@@
Let $\{X_{t}\}_{t=1,2,\ldots,T}$ follow model (11), with $f_{t}$ being a continuous, piecewise-linear signal and assume that the random sequence $\{\epsilon_{t}\}_{t=1,2,\ldots,T}$ is independent and identically distributed (i.i.d.) from the normal distribution with mean zero and variance one and also that (A1) and (A3) hold. Let $N$ and $r_{j},j=1,2,\ldots N$ be the number and location of the change-points, while $\hat{N}$ and $\hat{r}_{j},j=1,2,\ldots\hat{N}$ are their estimates, sorted in increasing order. In addition, $\Delta^{f}_{j}$, as defined in (13) is the magnitude of each change in the slope of $f_{t}$, and $\underline{f}_{T}$ and $\delta_{T}$ are defined in (12) and (14), respectively. Then there exist positive constants $C_{1},C_{2},C_{3},C_{4}$ independent of $T$ such that for $C_{1}\sqrt{\log T}\leq\zeta_{T}<C_{2}\delta_{T}^{3/2}\underline{f}_{T}$ and for sufficiently large $T$, we obtain
@@
The proof of Theorem 2 can be found in the supplementary material. Again, the lower bound of the probability is $1-\mathcal{O}(1/T)$ and the difficulty of the problem, $\delta_{T}^{3/2}\underline{f}_{T}$ is analogous to $\sqrt{\delta_{T}}\underline{f}_{T}$, which appears in (A2) for the case of piecewise-constant signals. For $\underline{f}_{T}\sim T^{1/2}$, the change-point detection accuracy is $O\left(T^{-1/3}\left(\log T\right)^{1/3}\right)$, which differs from the optimal rate $O(T^{-1/3})$ derived by [51] only by the logarithmic factor. This means that the change-point detection rate in DAIS is almost optimal.